1. Getty Markets has bonds outstanding that have a real return of 3.08 percent. The current rate of inflation is 3.1 percent. What is the nominal rate of return on these bonds?

6.28% Please show all work typed as I can't see much on mobile.
Answer #1
Nominal rate of return = (1+r)*(1+i) Where,
(1+R) = (1+0.0308)*(1+0.0310) r real rate of return = 3.08%
(1+R) =      1.0628 i Inflation rate = 3.10%
R =      1.0628 - 1 R Nominal rate of return = ?
R =      0.0628
So, Nominal rate of return is 6.28%